# The MORRIS Experiment: Magnetic Levitation as a New Probe of Non-Newtonian Gravity

### v1

@@ Conclusions
The total force imparted on our levitated particle due to the driving potential is given by Eq. 2. However, we pick out the signal at the fundamental frequency in Fourier space; this is given by the frequency at which the potential varies. This motivates us to decompose the full force into a constant component, $\mathcal{F}_{0}$, and an oscillatory component varying at the signal angular frequency $\omega_{s}$. We therefore approximate the total force at time $t$ and central displacement $r_{0}$ as
@@
This is the quantity governing our inferencing when searching for a fifth force, as given in Eq. 4.
Figure 4: Comparisons between the gravitational signal ($\alpha=0$) from the full calculation, Eq. 2, and our approximation, Eq. 7. This is shown in the time domain (top) and in the power spectral density (bottom) using Eq. 11. The oscillation amplitude $\mathcal{F}_{\mathrm{osc}}$ is given by the difference between the total force at the closest separation $a$ and farthest separation $a+2R$ for a single source mass. The approximation accurately captures the resonant peak, occurring at the signal frequency $f=f_{s}$ and which we use in our inferencing. The noise floor is assumed to be thermal and is computed in Eq. 5. We use our short-term setup, taking $r_{0}\approx 1.95\,\mathrm{cm}$.
We show the result of a simulation directly from Eq. 2 and from our approximation in Eq. 7 in Fig. 4 for the gravity-only scenario ($\alpha=0$). We visualize this using our short-term configuration (c.f. Table 1) in both the time domain and Fourier space; we use Eq. 11 for the latter. The PSD features the strongest peak at the fundamental frequency $f_{s}=\omega_{s}/(2\pi)$, with progressively weaker peaks at higher harmonics. Our approximation accurately captures the resonance peak, which is the part of the signal we plan to isolate in an experimental run. The value of the peak as predicted by our approximation is $\mathord{\sim}1.23$ times larger than that from the full simulation for any choice of the fifth-force parameters $\alpha$ and $\lambda$. Since both the gravitational and new physics signals scale by the same factor, we rely on our approximation and do not expect this difference to make an appreciable impact on our results. We also show the thermal noise floor, as computed in Eq. 5.
@@
We use the likelihood given in Eq. 6 to define the profile-likelihood ratio
$\mathop{{}\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0};\bm{p})}\equiv\frac{\mathop{{}\mathcal{L}\big{(}\alpha,\lambda_{0},\hat{\xi};\bm{p}\,\big{)}}}{\mathop{{}\mathcal{L}\big{(}\hat{\hat{\alpha}},\lambda_{0},\hat{\hat{\xi}};\bm{p}\,\big{)}}}\,,$ (13)
where single- and double-hats respectively denote the maximum likelihood estimators of the likelihood conditioned on a particular $(\alpha,\lambda)$ and of the unconditional likelihood. We fix to one value of the screening length $\lambda_{0}$ and perform a raster scan of the parameter space over different values of $\lambda_{0}$ [43].
@@
$t_{\alpha}\equiv-2\ln\left[\mathop{{}\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0});\bm{p}}\right]\,,$ (14)
where $\bm{p}\equiv\{p_{i}\}$ is our data vector. We use the TS to assess whether the null hypothesis, defined according to conditional likelihood, is a good description of the observed data compared to the alternative hypothesis, defined as per the maximized unconditional likelihood. We use a two-sided test statistic since $\alpha$ can take on both positive and negative values. A value of 00 for $t_{\alpha}$ indicates perfect agreement with the null hypothesis, whereas larger values indicate greater disagreement.
@@
To find $t_{\alpha}^{\mathrm{lim}}$, we must know the distributions of the test statistic for each choice of null hypothesis, $f(t_{\alpha}|\alpha,\lambda_{0})$. We generate these from a series of Monte Carlo (MC) simulations. For a large selection of null hypotheses, defined by the parameter space points $(\alpha,\lambda_{0})$, we generate $10^{6}$ pseudo-datasets distributed according to the likelihood given in Eq. 6 at our chosen central displacements. For each dataset, we then compute the value of $t_{\alpha}$ from Eq. 14, generating the desired TS distributions. For $\alpha$ values close to the ones we can exclude, we find that the distributions of $t_{\alpha}$ are well described by the asymptotic result of Ref. [39], with $t_{\alpha}$ following a $\chi^{2}$ distribution with one degree of freedom. This gives us a limiting TS value of $t_{\alpha}^{\mathrm{lim}}\approx 3.84$ for a $95\%$ CL limit.
@@
Figure 5: The $95\%$ confidence level projected limits on absolute values of $\alpha$ with screening length $\lambda$. We show the limits from our short-, medium-, and long-term configurations, summarized in Table 1. The dashed lines above unity $|\alpha|$ are thin, closed contours where negative values of $\alpha$ are allowed. The existing constraints and theoretical regions are discussed in the main text.
@@
We perform a sensitivity estimate by considering a single, oscillating source mass and by using an Asimov dataset that includes only the first data point [39]. For large non-centrality parameters $\kappa^{2}$, defined in Eq. 4, we can approximate the non-central $\chi^{2}$ distribution to be Gaussian. Its mean $\mu$ and variance $\sigma^{2}$ are equal to those of the original distribution, such that $\mu\simeq\kappa^{2}$ and $\sigma^{2}\simeq 4\kappa^{2}$. We allow for the non-centrality parameter to be shifted by the pull parameter $\xi$, governing the effects of experimental systematics and introduced in Eq. 6. Ignoring the penalty term on $\xi$, our approximate test statistic, formed from the log-likelihood ratio, is then
@@
$\begin{split}\mathcal{D}(\lambda,r)&\equiv\frac{(1+r/\lambda)\,e^{-r/\lambda}}{r^{2}}\,;\\ \Delta(\lambda)&\equiv\mathcal{D}(\lambda,a)-\mathcal{D}(\lambda,b)\,,\end{split}$ (18)
@@
Figure 6: The $95\%$ confidence level projected limits on the strength parameter $\alpha$ with screening length $\lambda$ from our complete statistical treatment and analytical approximations. We visualize the result for our medium-term configuration, summarized in Table 1. Below the minimum separation of $a_{\mathrm{min}}\approx 7\,\mathrm{mm}$, our approximation is based on the optimal sensitivity by setting the pull parameter $\xi$ to zero. Above this scale, systematics play a more dominant role, for which we set $\xi=-2\sigma_{\xi}$. The true limit lies between these cases, highlighted by the shaded region. The transparency of the analytical estimates approximately indicates where each one should be trusted.
@@
$\alpha_{\mathrm{lim}}(\lambda)\simeq\begin{cases}\pm 2\sigma_{\xi}/(1\mp 2\sigma_{\xi})\quad&\text{if}\hskip 6.45831pt\lambda\gg a_{\mathrm{min}}\\[4.30554pt] \pm\frac{\Delta(\infty)}{\Delta(\lambda)}\sqrt{t_{\mathrm{lim}}\frac{2S_{FF}^{\mathrm{noise}}}{\mathcal{F}_{g}^{2}T_{\mathrm{obs}}}}\quad&\text{if}\hskip 6.45831pt\lambda\lesssim a_{\mathrm{min}}\end{cases}\,.$ (22)
@@
×$\times\left[\left(\frac{40\,\mathrm{mg}}{m_{p}}\right)\left(\frac{1000\,\mathrm{h}}{T_{\mathrm{obs}}}\right)\left(\frac{10^{7}}{Q}\right)\left(\frac{T}{1\,\mathrm{K}}\right)\left(\frac{f_{s}}{35\,\mathrm{Hz}}\right)\right]^{1/2}\,,$
where we have fiducialized to our medium-term configuration (c.f. Table 1). We see that the optimal parameters to consider are the minimal separation distance and the mass of the source.
@@
The experimental values shown in Table 1 are chosen to maximize our sensitivity while minimizing the experimental difficulty in their implementation. We justify our proposed parameters below.
@@
In the short-term setup, our precision is mainly limited by our temperature readout, the frequency stability of the drive, and the precision with which we can determine the sensing mass. The effect of the uncertainty in the mass can be minimized by the rejection of systematic errors. The thermal and frequency stabilities are sources of non-systematic noise that dominate the uncertainty in the pull parameter. Improvements on the thermal stability can be readily made to reduce this noise for the medium- and long-term setups by increasing the resolution of the thermometers used and employing a PID scheme to stabilize the temperature of the experiment. The frequency stability can be increased by using higher resolution optical encoders and dedicated driving electronics for use at low frequencies. At high $Q$-factors, this frequency needs to be stable to better than $f/Q$. For the short- and medium-term setups, this stability can be readily achieved through commercially available optical encoders. For the long-term setup, detected electronics must be used to create a stable Phase-Locked Loop.

### v2

@@ Conclusions
The total force imparted on our levitated particle due to the driving potential is given by the sum of the individual forces from each source mass. However, we pick out the signal at the fundamental frequency in Fourier space; this is given by the frequency at which the potential varies. This motivates us to decompose the full force into a constant component, $\mathcal{F}_{0}$, and an oscillatory component varying at the signal angular frequency $\omega_{s}$. We therefore approximate the total force at time $t$ and central displacement $r_{0}$ as
@@
This is the quantity governing our inferencing when searching for a fifth force.
Figure 4: Comparisons between the gravitational signal ($\alpha=0$) from the full calculation and our approximation, Eq. 7. This is shown in the time domain (top) and in the power spectral density (bottom) using Eq. 11. The oscillation amplitude $\mathcal{F}_{\mathrm{osc}}$ is given by the difference between the total force at the closest separation $a$ and farthest separation $a+2R$ for a single source mass. The approximation accurately captures the resonant peak, occurring at the signal frequency $f=f_{s}$ and which we use in our inferencing. The noise floor is assumed to be thermal and is computed in the main text. We use our short-term setup, taking $r_{0}\approx 1.95\,\mathrm{cm}$.
We show the result of a simulation using the full force treatment and the approximation given by Eq. 7 in Fig. 4 for the gravity-only scenario ($\alpha=0$). We visualize this using our short-term configuration in both the time domain and Fourier space; we use Eq. 11 for the latter. The PSD features the strongest peak at the fundamental frequency $f_{s}=\omega_{s}/(2\pi)$, with progressively weaker peaks at higher harmonics. Our approximation accurately captures the resonance peak, which is the part of the signal we plan to isolate in an experimental run. The value of the peak as predicted by our approximation is $\mathord{\sim}1.23$ times larger than that from the full simulation for any choice of the fifth-force parameters $\alpha$ and $\lambda$. Since both the gravitational and new physics signals scale by the same factor, we rely on our approximation and do not expect this difference to make an appreciable impact on our results. We also show the thermal noise floor, as computed in the main text.
@@
Using the likelihood given in the main text, $\mathcal{L}$, we define the profile-likelihood ratio
$\mathop{{}\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0};\boldsymbol{p})}\equiv\frac{\mathop{{}\mathcal{L}\big(\alpha,\lambda_{0},\hat{\xi};\boldsymbol{p}\,\big)}}{\mathop{{}\mathcal{L}\big(\hat{\hat{\alpha}},\lambda_{0},\hat{\hat{\xi}};\boldsymbol{p}\,\big)}}\,$, (13)
where $\boldsymbol{p}\equiv\{p_{i}\}$ is our data vector. The single- and double-hats respectively denote the maximum likelihood estimators of the likelihood conditioned on a particular $(\alpha,\lambda)$ and of the unconditional likelihood. We fix to one value of the screening length $\lambda_{0}$ and perform a raster scan of the parameter space over different values of $\lambda_{0}$ [6].
@@
$t_{\alpha}\equiv-2\ln\left[\mathop{{}\lambda_{\mathrm{PLR}}(\alpha,\lambda=\lambda_{0});\boldsymbol{p}}\right]\,.$ (14)
We use the TS to assess whether the null hypothesis, defined according to conditional likelihood, is a good description of the observed data compared to the alternative hypothesis, defined as per the maximized unconditional likelihood. We use a two-sided test statistic since $\alpha$ can take on both positive and negative values. A value of 0 for $t_{\alpha}$ indicates perfect agreement with the null hypothesis, whereas larger values indicate greater disagreement.
@@
To find $t_{\alpha}^{\mathrm{lim}}$, we must know the distributions of the test statistic for each choice of null hypothesis, $f(t_{\alpha}|\alpha,\lambda_{0})$. We generate these from a series of Monte Carlo (MC) simulations. For a large selection of null hypotheses, defined by the parameter space points $(\alpha,\lambda_{0})$, we generate $10^{6}$ pseudo-datasets distributed according to our likelihood function at our chosen central displacements. For each dataset, we then compute the value of $t_{\alpha}$ from Eq. 14, generating the desired TS distributions. For $\alpha$ values close to the ones we can exclude, we find that the distributions of $t_{\alpha}$ are well described by the asymptotic result of Ref. [10], with $t_{\alpha}$ following a $\chi^{2}$ distribution with one degree of freedom. This gives us a limiting TS value of $t_{\alpha}^{\mathrm{lim}}\approx 3.84$ for a $95\%$ CL limit.
@@
Figure 5: The $95\%$ confidence level projected limits on absolute values of $\alpha$ with screening length $\lambda$. We show the limits from our short-, medium-, and long-term configurations. The dashed lines above unity $|\alpha|$ are thin, closed contours where negative values of $\alpha$ are allowed. The existing constraints and theoretical regions are discussed in the main text.
@@
We perform a sensitivity estimate by considering a single, oscillating source mass and by using an Asimov dataset that includes only the first data point [10]. For large non-centrality parameters $\kappa^{2}$, defined in Eq. 12, we can approximate the non-central $\chi^{2}$ distribution to be Gaussian. Its mean $\mu$ and variance $\sigma^{2}$ are equal to those of the original distribution, such that $\mu\simeq\kappa^{2}$ and $\sigma^{2}\simeq 4\kappa^{2}$. We allow for the non-centrality parameter to be shifted by the pull parameter $\xi$, governing the effects of experimental systematics. Ignoring the penalty term on $\xi$, our approximate test statistic, formed from the log-likelihood ratio, is then
@@
[…]
@@
Figure 6: The $95\%$ confidence level projected limits on the strength parameter $\alpha$ with screening length $\lambda$ from our complete statistical treatment and analytical approximations. We visualize the result for our medium-term configuration. Below the minimum separation of $a_{\mathrm{min}}\approx 7\,\mathrm{mm}$, our approximation is based on the optimal sensitivity by setting the pull parameter $\xi$ to zero. Above this scale, systematics play a more dominant role, for which we set $\xi=-2\sigma_{\xi}$. The true limit lies between these cases, highlighted by the shaded region. The transparency of the analytical estimates approximately indicates where each one should be trusted.
@@
$\hskip 0.0pt\alpha_{\mathrm{lim}}(\lambda)\simeq\begin{cases}\pm 2\sigma_{\xi}/(1\mp 2\sigma_{\xi})\quad&\text{if}\hskip 6.45831pt\lambda\gg a_{\mathrm{min}}\\[4.30554pt] \pm\frac{\Delta(\infty)}{\Delta(\lambda)}\sqrt{t_{\mathrm{lim}}\frac{2S_{FF}^{\mathrm{noise}}}{\mathcal{F}_{g}^{2}T_{\mathrm{obs}}}}\quad&\text{if}\hskip 6.45831pt\lambda\lesssim a_{\mathrm{min}}\end{cases}\,.$ (22)
@@
$\times\left[\left(\frac{40\,\mathrm{mg}}{m_{p}}\right)\left(\frac{1000\,\mathrm{h}}{T_{\mathrm{obs}}}\right)\left(\frac{10^{7}}{Q}\right)\left(\frac{T}{1\,\mathrm{K}}\right)\left(\frac{f_{s}}{35\,\mathrm{Hz}}\right)\right]^{1/2}\,$,
where we have fiducialized to our medium-term configuration. We see that the optimal parameters to consider are the minimal separation distance and the mass of the source.
@@
We chose our experimental values to maximize our sensitivity while minimizing the experimental difficulty in their implementation. We justify our proposed parameters below.
@@
In the short-term setup, our precision is mainly limited by our temperature readout, the frequency stability of the drive, and the precision with which we can determine the sensing mass. Based on a selection of commercially available methods to measure and control these values, the dominant uncertainty is expected to be the temperature. Thermocouples are wildly used for temperature sensing, but suffer from low signal-to-noise ratios at deep cryogenic temperatures, from which we estimate a $10\%$ dominant uncertainty in the short term. Use of commercial resistive and magnetic susceptibility thermometers would reduce this to be subdominant, in the order of $10\,\mathrm{\mu K}$ on a proposed temperature of $100\,\mathrm{mK}$, for an error of $0.01\%$.
Commercial optical encoders typically have anywhere from $500$ to $100,000$ lines per revolution. On a frequency of $35\,\mathrm{Hz}$, our disk spins at roughly $10\,\mathrm{Hz}$, or a $100\mathrm{\mu Hz}$ frequency resolution using a $100,000$ line encoder, leading to an error of $0.01\%$, which is subdominant in both short and medium term set-up. In the long term set-up, where a second particle is driven, optical encoders cannot be used, so we conservatively estimate a frequency uncertainty of $0.1\%$, which driven primarily by the phase noise in the system and the performance of typical PLL loops. This is the dominant uncertainty in the long term configuration.
To determine the source masses, commercial mass balance scales provide resolutions of up to $0.1\,\mathrm{\mu g}$ on a mass of $10\,\mathrm{g}$, resulting in a fractional uncertainty of $10^{-6}\%$. However, for the medium term, our total mass is larger than this total allowed mass. We therefore require less sensitive scales, so we conservatively estimate a dominant uncertainty of $1\%$ in the medium-term configuration arising from the mass uncertainty.
We note that the mass uncertainty is an important issue in any small-scale gravity measurement, and therefore also for a fifth-force search using these techniques. In a similar fashion to how we turn our positional uncertainty into a systematic error via our pull parameter, which does not play an appreciable role when the slope in our estimator is manifest, we eventually aim to use several different mass wheels with slightly different source masses. When the masses are measured via the same technique (either through balance scales or though SEM volume measures possible at this mass scale), this moves our mass uncertainty to a systematic error. This is also largely eliminated through the use of the slope.
The thermal and frequency stabilities are sources of non-systematic noise that dominate the uncertainty in the pull parameter. Improvements on the thermal stability can be readily made to reduce this noise for the medium- and long-term setups by increasing the resolution of the thermometers used and employing a PID scheme to stabilize the temperature of the experiment. The frequency stability can be increased by using higher resolution optical encoders and dedicated driving electronics for use at low frequencies. At high $Q$-factors, this frequency needs to be stable to better than $f/Q$. For the short- and medium-term setups, this stability can be readily achieved through commercially available optical encoders. For the long-term setup, dedicated electronics must be used to create a stable Phase-Locked Loop.